# The impact of cosmic filaments on the abundance of satellite galaxies

### v1

@@ 1 Introduction
In this study, we utilize one of the state-of-the-art cosmological hydrodynamical simulations, IllustrisTNG (F. Marinacci et al., 2018; J. P. Naiman et al., 2018; D. Nelson et al., 2018, 2019; A. Pillepich et al., 2018a, b; V. Springel et al., 2018), which allows us to investigate the impact of filaments on satellite abundance with direct insights into the physical origins of environmental effects. The paper is organized as follows. In Section 2, we describe the details of the IllustrisTNG simulations, the filament identification algorithm, and the definition of environment. In Section 3, we present the impact of filamentary environment on satellite abundance, including comparisons with observations, the impact of magnitude distributions, and the influence of halo mass distributions. The effects of tracer choice are investigated in Section 4. Finally, we summarize our findings and conclude in Section 5.
2 Methods
2.1 Simulation
Table 1: Number of central galaxies in bins of $r$-band absolute magnitude ($M_{r,{\rm cen}}$) and different cosmic web environments.
Bins $M_{r,{\rm cen}}$ range Total Knot Galaxy tracers Dark matter tracers
@@ 2.1 Simulation
Halos and subhalos in the simulation are identified using the friends-of-friends (FoF, M. Davis et al., 1985) and SUBFIND (V. Springel et al., 2001; K. Dolag et al., 2009) algorithms, respectively. We define the host halo mass as $M_{\mathrm{200c}}$, the total mass enclosed within the virial radius $R_{\mathrm{200c}}$, within which the mean density is 200 times the cosmic critical density. The galaxy sample is based on the SUBFIND subhalo catalogue in the TNG100-1 run. We exclude subhalos of non-cosmological origin (e.g., fragments or clumps formed through baryonic processes within existing galaxies, which are marked with SubhaloFlag=0 in the SUBFIND subhalo catalog; see section 5.2 in D. Nelson et al. 2019) as well as low-mass subhalos containing fewer than 100 star particles, which corresponds to satellite galaxies with stellar masses of $M_{\mathrm{star}}\sim 10^{8}~M_{\odot}$ or $r$-band magnitudes $M_{r}\sim-14.6$ (with a typical scatter of $\Delta M_{r}=0.6$). In this study, the stellar mass and $r$-band magnitude of each SUBFIND structure are computed using star particles within a 3D radial aperture of 30 physical kpc, consistent with the SDSS Petrosian aperture. To compute the satellite abundance of each FoF halo, we construct corresponding satellite samples. The central galaxy of each FoF halo is defined as the most bound SUBFIND subhalo in the catalogue, while satellite galaxies are defined as the remaining subhalos that meet the above criteria and reside within the virial radius of their host halo.
@@
For a more direct comparison with the observational results in G15, we consider the SDSS $r$-band absolute magnitude of galaxies, $M_{r}$, in this study. We use $M_{r}$ data from D. Nelson et al. (2018), which are based on a model that accounts for dust attenuation due to the distribution of neutral gas and metals (specifically, ‘Model C’ in D. Nelson et al. 2018). We include all central galaxies with magnitudes in the range $23.5\leq M_{r,{\rm cen}}\leq-19.5$, and divide them into four magnitude bins. The bright-end limit of $M_{r,{\rm cen}}=-23.5$ is motivated by G15, while the faint-end limit of $M_{r,{\rm cen}}=-19.5$ is chosen to maximize the sample size within the range of reliable numerical resolution. The number of central galaxies in each bin is listed in the third column of Table 1.
2.2 Filament identification
@@
Figure 1: Filaments in a $30$ Mpc-thick slice of the TNG100-1 simulation, identified by the DisPerSE algorithm using galaxies with stellar masses $M_{\rm star}\geq 10^{9}~M_{\odot}$ as tracers. The grayscale background shows the dark matter density field. White circles mark the $3\times R_{\mathrm{200c}}$ vicinities of galaxy clusters with $M_{\mathrm{200c}}\geq 10^{13.5}~M_{\odot}$, which are defined as the knot environment. Red dots indicate the centers of galaxies residing in knots, which are excluded from our analysis. Black solid line segments show the filament spines defined by DisPerSE. Orange dots mark the centers of galaxies located within cylinders of radius of $R_{\rm filament}=1$ Mpc around the filament spines (indicated by the orange shaded regions), and defined as filament galaxies. The remaining galaxies, shown as blue dots, are field galaxies. Visually, despite some minor discrepancies, the filament spines overall capture the massive filamentary structures in the underlying dark matter density field.
@@
Following observations, we use galaxies as tracers for filament identification. Similar to D. Galárraga-Espinosa et al. (2024), we select galaxies according to observational limits, balancing the identification of real structures with the suppression of spurious ones. Specifically, we trace filaments using galaxies (including both central and satellite galaxies) with stellar masses $M_{\mathrm{star}}\geq 10^{9}~M_{\odot}$, apply smoothing on both the estimated density field and the identified filaments (each smoothed once), and select filaments with a persistence ratio above the threshold of $2\sigma$. Note that we have tested a range of persistence ratio thresholds (from $1\sigma$ to $6\sigma$), with and without smoothing procedures applied, and confirmed that the conclusions presented in the following sections are not sensitive to these choices.
The identified filament spines in a slice of the simulation box are shown as black solid lines in Figure 1. Overall, the filament spines capture the massive filamentary structures of the underlying dark matter density field (shown as the grayscale background), despite minor mismatches in smaller features. These discrepancies arise primarily from the use of a 2D projection and from galaxies being a sparse tracer of the density field. The impact of tracers will be examined in Section 4.
2.3 Definition of environment
After identifying the filamentary spines, we further classify central galaxies into ‘filament’ and ‘field’ environments. To do this, we first exclude the galaxies residing in clusters (i.e., the ‘knot’ environment), as these environments have distinct effects on galaxy properties compared to filaments and fields. Specifically, central galaxies located within $3R_{\mathrm{200c}}$ of clusters with $M_{\mathrm{200c}}\geq 10^{13.5}~M_{\odot}$ are classified as ‘knot’ environment galaxies (i.e., galaxies marked by red dots in Figure 1) and are excluded from our galaxy samples. The numbers of ‘knot’ galaxies in different $M_{r,{\rm cen}}$ bins are listed in the fourth column of Table 1.
@@ 2.3 Definition of environment
3 Satellite luminosity functions
3.1 Environmental dependence
@@
where $N_{{\rm sat},i}(M_{\rm diff})$ is the number of satellites associated with the $i$-th central galaxy whose magnitude differences fall within the bin $M_{\rm diff}-\Delta M_{\rm diff}/2,M_{\rm diff}+\Delta M_{\rm diff}/2)$, and $N_{\rm cen}$ is the total number of central galaxies considered. We also estimate the uncertainty in the satellite LF, shown as shaded regions, based on Poisson errors, i.e.,
@@
Figure 3: Top: Distributions of central galaxy magnitudes ($M_{r,{\rm cen}}$). From left to right, we show the distributions of the original samples, the samples after resampling $M_{r,{\rm cen}}$, and the samples after resampling both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$. In each panel, as in other figures in this study, the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins are plotted in red, green, and yellow, respectively. Central galaxies in filaments and the field are distinguished using solid and dashed lines, respectively. Bottom: Same as the top panels, but showing the distributions of the host halo virial masses ($M_{\rm 200c}$) for central galaxies. After resampling, the filament and field samples exhibit similar distributions in both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$. See the main text for details of the resampling procedure.
@@
Original Resampling $M_{r,{\rm cen}}$ Resampling $M_{r,{\rm cen}}$ and $M_{\mathrm{200c}}$
@@
3.2 Impact of magnitude distributions
@@
To further quantify the impact of magnitude distribution on the environmental difference observed in Section 3.1, we resample the central galaxies based on their $M_{r,{\rm cen}}$, ensuring that the resampled samples have similar distributions in both environments. Specifically, for each pair of filament and field histograms, we divide them into 20 narrower bins. [^1] We have tested other choices for the number of narrow bins (e.g., 10, 40) and confirmed that the results remain consistent. A similar test is also performed for the resampling of $M_{\rm 200c}$ (Section 3.3). In each narrow bin, central galaxies from the environment with a higher number count are randomly resampled to match the number of galaxies in the other environment. The $M_{r,{\rm cen}}$ distributions of the resampled central galaxy samples are shown in the upper-middle panel of Figure 3. We can clearly see that, after resampling, the filament and field histograms are closely matched. We then re-examine the environmental differences in satellite LFs using these resampled central galaxies.
The satellite LFs computed from resampled central galaxies across different magnitude bins are plotted in the middle column of Figure 4. Compared to the original results shown in the left column, the resampled LFs exhibit larger Poisson errors due to reduced number counts. However, the ratio panels reveal that the environmental differences remain largely unchanged, with average ratios of $3.07$, $2.54$, and $1.92$ for the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins, respectively (see the fourth column in Table 2). Compared to the original ratios, the resampled values decrease by less than $17\%$. This suggests that the difference in magnitude distributions accounts for only a small part of the observed environmental effect, and that the environmental difference reported in Section 3.1 likely arises from other factors.
3.3 Impact of halo mass distributions
@@
To quantify the impact of halo mass distributions, we follow the procedure described in Section 3.2. Starting from the resampled central galaxy samples with matched $M_{r,{\rm cen}}$ distributions, we further divide the halo mass distributions (shown in the bottom-middle panel of Figure 3) into 20 narrower bins. Within each narrow bin, we randomly resample central galaxies in the two environments to ensure equal numbers. This procedure yields central galaxy samples with very similar distributions in both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$, as shown in the right column of Figure 3.
@@ 3.4 Discussions
4 Tracer Effects
Figure 5: Similar to Figure 1, but for filaments identified using dark matter density as tracers. The spines of dark matter-traced filaments are shown as purple solid lines, while the galaxy-traced filament spines from Figure 1 are overlaid using black dotted lines for comparison. Filaments identified by these two tracers are qualitatively similar, both following the underlying filamentary matter distribution, but they differ in detailed structures. We investigate these differences and their impact in Section 4.
@@ 4 Tracer Effects
4.1 Dark matter-traced filaments
It has been shown that the choice of tracers can affect the detailed properties of filaments identified by DisPerSE (see e.g., C. Laigle et al., 2018; D. Zakharova et al., 2023; Y. M. Bahe & P. Jablonka, 2025). Using cosmological simulations, these studies demonstrate that filament properties – such as spatial distributions, lengths, and connectivity – depend on whether dark matter particles or galaxies are used as tracers. This discrepancy arises because galaxies, distributed sparsely and discontinuously, are biased tracers of the underlying matter distribution, whereas dark matter more faithfully traces the total mass density field. In observations, we are limited to using galaxies as tracers, as was done in Section 3. In this section, we take advantage of simulations to explore the use of dark matter as tracers and investigate how the tracer choice affects the observed environmental dependence of satellite galaxy abundance.
For dark matter tracers, following Y. M. Bahe & P. Jablonka (2025), we use a smoothed density field instead of randomly sampled dark matter particles. This choice has minimal impact on the resulting filaments but yields more consistent identification with fewer artifacts at comparable computational expense. Specifically, we generate a gridded 3D density field with a cell size of 83.3 kpc, apply Gaussian smoothing with a scale of 500 kpc, and then downsample the field to a coarser grid with 250 kpc cells to reduce computational cost. Note that, since the numbers of tracer particles differ between different choices, it is usually not possible to use identical parameter sets in the DisPerSE finder. Following D. Zakharova et al. (2023), and to enable a more direct and fair comparison with galaxy-traced filaments, we adopt a higher persistence ratio threshold of $5.1\sigma$ (compared to $2\sigma$ for galaxy tracers), ensuring that the total filament lengths are matched between the two cases (i.e., $L_{\rm tot,fiducial}=7.84\times 10^{3}$ Mpc).

### v2

@@ I Introduction
In this study, we utilize one of the state-of-the-art cosmological hydrodynamical simulations, IllustrisTNG (F. Marinacci et al., 2018; J. P. Naiman et al., 2018; D. Nelson et al., 2018, 2019; A. Pillepich et al., 2018a, b; V. Springel et al., 2018), which allows us to investigate the impact of filaments on satellite abundance with direct insights into the physical origins of environmental effects. The paper is organized as follows. In Section II, we describe the details of the IllustrisTNG simulations, the filament identification algorithm, and the definition of environment. In Section III, we present the impact of filamentary environment on satellite abundance, including comparisons with observations, the impact of magnitude distributions, and the influence of halo mass distributions. The effects of tracer choice are investigated in Section IV. Finally, we summarize our findings and conclude in Section V.
II Methods
II.1 Simulation
Table 1: Number of central galaxies (and their associated satellites) in bins of $r$-band absolute magnitude ($M_{r,{\rm cen}}$) and in different cosmic web environments. Each bin includes all centrals with $M_{r,{\rm cen}}$ within $\pm 0.5$ mag of the bin center. The numbers in parentheses give the total number of satellites (with more than 100 stellar particles and located within the $R_{\rm 200c}$ of their host) associated with the centrals in each bin. The listed stellar mass ranges correspond to the 5th–95th percentile of all central galaxies, irrespective of environment. Note that the centrals in each bin are selected by $M_{r,{\rm cen}}$, so their stellar mass ranges slightly overlap. For both tracers, the total number of central galaxies in the combined filament and field environments is the same in each magnitude bin, owing to our environment definition (see Section II.3).
Bins $M_{\rm star}$ range Total Knot Galaxy tracers Dark matter tracers
@@ II.1 Simulation
Halos and subhalos in the simulation are identified using the friends-of-friends (FoF, M. Davis et al., 1985) and SUBFIND (V. Springel et al., 2001; K. Dolag et al., 2009) algorithms, respectively. We define the host halo mass as $M_{\rm 200c}$, the total mass enclosed within the virial radius $R_{\rm 200c}$, within which the mean density is 200 times the cosmic critical density. The galaxy sample is based on the SUBFIND subhalo catalogue in the TNG100-1 run. We exclude subhalos of non-cosmological origin (e.g., fragments or clumps formed through baryonic processes within existing galaxies, which are marked with SubhaloFlag=0 in the SUBFIND subhalo catalog; see section 5.2 in D. Nelson et al. 2019) as well as low-mass subhalos containing fewer than 100 star particles, which corresponds to satellite galaxies with stellar masses of $M_{\rm star}\sim 10^{8}~M_{\odot}$ or $r$-band magnitudes $M_{r}\sim-14.6$ (with a typical scatter of $\Delta M_{r}=0.6$). In this study, the stellar mass and $r$-band magnitude of each SUBFIND structure are computed using star particles within a 3D radial aperture of 30 physical kpc, consistent with the SDSS Petrosian aperture. To compute the satellite abundance of each FoF halo, we construct corresponding satellite samples. The central galaxy of each FoF halo is defined as the most bound SUBFIND subhalo in the catalogue, while satellite galaxies are defined as the remaining subhalos that meet the above criteria and reside within the virial radius of their host halo.
@@
For a more direct comparison with the observational results in G15, we consider the SDSS $r$-band absolute magnitude of galaxies, $M_{r}$, in this study. We use $M_{r}$ data from D. Nelson et al. (2018), which are based on a model that accounts for internal dust attenuation arising from the distribution of neutral gas and metals within galaxies (specifically, ‘Model C’ in D. Nelson et al. 2018).  [^1] We have tested and confirmed that our main conclusions are not sensitive to the dust model and remain consistent when using dust-free $M_{r}$. We include all central galaxies with magnitudes in the range $23.5\leq M_{r,{\rm cen}}\leq-19.5$, and divide them into four magnitude bins. The bright-end limit of $M_{r,{\rm cen}}=-23.5$ is motivated by G15, while the faint-end limit of $M_{r,{\rm cen}}=-19.5$ is chosen to maximize the sample size within the range of reliable numerical resolution. The third column of Table 1 lists the total number of central galaxies in each bin (with the corresponding satellite counts given in parentheses), while the second column provides the typical stellar mass range of centrals.
II.2 Filament identification
@@
Figure 1: Filaments in a $30$ Mpc-thick slice of the TNG100-1 simulation, identified by the DisPerSE algorithm using galaxies with stellar masses $M_{\rm star}\geq 10^{9}~M_{\odot}$ as tracers. The grayscale background shows the dark matter density field. White circles mark the $3\times R_{\rm 200c}$ vicinities of galaxy clusters with $M_{\rm 200c}\geq 10^{13.5}~M_{\odot}$, which are defined as the knot environment. Red dots indicate the centers of galaxies residing in knots, which are excluded from our analysis. Black solid line segments show the filament spines defined by DisPerSE. Orange dots mark the centers of galaxies located within cylinders of radius of $R_{\rm filament}=1$ Mpc around the filament spines (indicated by the orange shaded regions), and defined as filament galaxies. The remaining galaxies, shown as blue dots, are field galaxies. Visually, despite some minor discrepancies, the filament spines overall capture the massive filamentary structures in the underlying dark matter density field.
@@
Following observations, we use galaxies as tracers for filament identification. Similar to D. Galárraga-Espinosa et al. (2024), we select galaxies according to observational limits, balancing the identification of real structures with the suppression of spurious ones. Specifically, we trace filaments using galaxies (including both central and satellite galaxies) with stellar masses $M_{\rm star}\geq 10^{9}~M_{\odot}$, apply smoothing on both the estimated density field and the identified filaments (each smoothed once), and select filaments with a persistence ratio above the threshold of $2\sigma$. Note that we have tested a range of persistence ratio thresholds (from $1\sigma$ to $6\sigma$, see Appendix A for details), with and without smoothing procedures applied, and confirmed that the conclusions presented in the following sections are not sensitive to these choices.
The identified filament spines in a slice of the simulation box are shown as black solid lines in Figure 1. Overall, the filament spines capture the massive filamentary structures of the underlying dark matter density field (shown as the grayscale background), despite minor mismatches in smaller features. These discrepancies arise primarily from the use of a 2D projection and from galaxies being a sparse tracer of the density field. The impact of tracers will be examined in Section IV.
II.3 Definition of environment
After identifying the filamentary spines, we further classify central galaxies into ‘filament’ and ‘field’ environments. To do this, we first exclude the galaxies residing in clusters (i.e., the ‘knot’ environment), as these environments have distinct effects on galaxy properties compared to filaments and fields. Specifically, central galaxies located within $3R_{\rm 200c}$ of clusters with $M_{\rm 200c}\geq 10^{13.5}~M_{\odot}$ are classified as ‘knot’ environment galaxies (i.e., galaxies marked by red dots in Figure 1) and are excluded from our galaxy samples. The numbers of ‘knot’ galaxies in different $M_{r,{\rm cen}}$ bins are listed in the fourth column of Table 1.
@@ II.3 Definition of environment
III Satellite luminosity functions
III.1 Environmental dependence
@@
where $N_{{\rm sat},i}(M_{\rm diff})$ is the number of satellites associated with the $i$-th central galaxy whose magnitude differences fall within the bin $M_{\rm diff}-\Delta M_{\rm diff}/2,M_{\rm diff}+\Delta M_{\rm diff}/2)$, $\Delta M_{\rm diff}$ denotes the bin width of the magnitude difference, and $N_{\rm cen}$ is the total number of central galaxies considered. We also estimate the uncertainty in the satellite LF, shown as shaded regions, based on Poisson errors, i.e.,
@@
Figure 3: Top: Distributions of central galaxy magnitudes ($M_{r,{\rm cen}}$). From left to right, we show the distributions of the original samples, the samples after resampling $M_{r,{\rm cen}}$, and the samples after resampling both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$. In each panel, as in other figures in this study, the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins are plotted in red, green, and yellow, respectively. Central galaxies in filaments and the field are distinguished using solid and dashed lines, respectively. For each $M_{r,{\rm cen}}$ bin, the Kolmogorov–Smirnov (KS) test $p$-value is shown in the panel, quantifying the similarity between the filament and field $M_{r,{\rm cen}}$ distributions. Bottom: Same as the top panels, but showing the distributions of the host halo virial masses ($M_{\rm 200c}$) for central galaxies. After resampling, the filament and field samples exhibit similar distributions in both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$. See the main text for details of the resampling procedure.
@@
Original Resampling $M_{r,{\rm cen}}$ Resampling $M_{r,{\rm cen}}$ and $M_{\rm 200c}$
@@
III.2 Impact of magnitude distributions
@@
To further quantify the impact of magnitude distribution on the environmental difference observed in Section III.1, we resample the central galaxies based on their $M_{r,{\rm cen}}$, ensuring that the resampled samples have similar distributions in both environments. Specifically, for each pair of filament and field histograms, we divide them into 20 narrower bins. [^2] We have tested other choices for the number of narrow bins (e.g., 10, 40) and confirmed that the results remain consistent. A similar test is also performed for the resampling of $M_{\rm 200c}$ (Section III.3). In each narrow bin, central galaxies from the environment with a higher number count are randomly resampled to match the number of galaxies in the other environment. The $M_{r,{\rm cen}}$ distributions of the resampled central galaxy samples are shown in the upper-middle panel of Figure 3. We can clearly see that, after resampling, the filament and field histograms are closely matched, as further confirmed by the Kolmogorov–Smirnov (KS) tests shown in Figure 3 (i.e., a higher $p$-value indicating that the two distributions are statistically indistinguishable). We then re-examine the environmental differences in satellite LFs using these resampled central galaxies.
The satellite LFs computed from resampled central galaxies across different magnitude bins are plotted in the middle column of Figure 4. Compared to the original results shown in the left column, the resampled LFs exhibit larger Poisson errors due to reduced number counts. However, the ratio panels reveal that the environmental differences remain largely unchanged, with average ratios of $3.07$, $2.54$, and $1.92$ for the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins, respectively (see the fourth column in Table 2). Compared to the original ratios, the resampled values decrease by less than $17\%$. This suggests that the difference in magnitude distributions accounts for only a small part of the observed environmental effect, and that the environmental difference reported in Section III.1 likely arises from other factors.
III.3 Impact of halo mass distributions
@@
To quantify the impact of halo mass distributions, we follow the procedure described in Section III.2. Starting from the resampled central galaxy samples with matched $M_{r,{\rm cen}}$ distributions, we further divide the halo mass distributions (shown in the bottom-middle panel of Figure 3) into 20 narrower bins. Within each narrow bin, we randomly resample central galaxies in the two environments to ensure equal numbers. This procedure yields central galaxy samples with very similar distributions in both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$, as shown in the right column of Figure 3.
@@ III.4 Discussions
IV Tracer Effects
Figure 5: Similar to Figure 1, but for filaments identified using dark matter density as tracers. The spines of dark matter-traced filaments are shown as purple solid lines, while the galaxy-traced filament spines from Figure 1 are overlaid using black dotted lines for comparison. Filaments identified by these two tracers are qualitatively similar, both following the underlying filamentary matter distribution, but they differ in detailed structures. We investigate these differences and their impact in Section IV.
@@ IV Tracer Effects
IV.1 Dark matter-traced filaments
It has been shown that the choice of tracers can affect the detailed properties of filaments identified by DisPerSE (see e.g., C. Laigle et al., 2018; D. Zakharova et al., 2023; Y. M. Bahé & P. Jablonka, 2025). Using cosmological simulations, these studies demonstrate that filament properties – such as spatial distributions, lengths, and connectivity – depend on whether dark matter particles or galaxies are used as tracers. This discrepancy arises because galaxies, distributed sparsely and discontinuously, are biased tracers of the underlying matter distribution, whereas dark matter more faithfully traces the total mass density field. In observations, we are limited to using galaxies as tracers, as was done in Section III. In this section, we take advantage of simulations to explore the use of dark matter as tracers and investigate how the tracer choice affects the observed environmental dependence of satellite galaxy abundance.
For dark matter tracers, following Y. M. Bahé & P. Jablonka (2025), we use a smoothed density field instead of randomly sampled dark matter particles. This choice has minimal impact on the resulting filaments but yields more consistent identification with fewer artifacts at comparable computational expense. Specifically, we generate a gridded 3D density field with a cell size of 83.3 kpc, apply Gaussian smoothing with a scale of 500 kpc, and then downsample the field to a coarser grid with 250 kpc cells to reduce computational cost. Note that, since the numbers of tracer particles differ between different choices, it is usually not possible to use identical parameter sets in the DisPerSE finder. Following D. Zakharova et al. (2023), and to enable a more direct and fair comparison with galaxy-traced filaments, we adopt a higher persistence ratio threshold of $5.1\sigma$ (compared to $2\sigma$ for galaxy tracers), ensuring that the total filament lengths are matched between the two cases (i.e., $L_{\rm tot,fiducial}=7.84\times 10^{3}$ Mpc). Alternatively, we have tested selecting the persistence ratio to match the number of galaxies in filaments to that of the galaxy-traced case, and found that this does not affect our main conclusions.